# NM-FlowGAN: Modeling sRGB Noise with a Hybrid Approach based on Normalizing Flows and Generative Adversarial Networks

### v2

@@ Abstract
Modeling and synthesizing real sRGB noise is crucial for various low-level vision tasks, such as building datasets for training image denoising systems. The distribution of real sRGB noise is highly complex and affected by a multitude of factors, making its accurate modeling extremely challenging. Therefore, recent studies have proposed methods that employ data-driven generative models, such as generative adversarial networks (GAN) and Normalizing Flows. These studies achieve more accurate modeling of sRGB noise compared to traditional noise modeling methods. However, there are performance limitations due to the inherent characteristics of each generative model. To address this issue, we propose NM-FlowGAN, a hybrid approach that exploits the strengths of both GAN and Normalizing Flows. We simultaneously employ a pixel-wise noise modeling network based on Normalizing Flows, and spatial correlation modeling networks based on GAN. In our experiments, our NM-FlowGAN outperforms other baselines on the sRGB noise synthesis task. Moreover, the denoising neural network, trained with synthesized image pairs from our model, also shows superior performance compared to other baselines. Our code is available at: https://github.com/YoungJooHan/NM-FlowGAN.
Keywords:
Noise Modeling Noise Synthesizing Image Denoising
1 Introduction
Real-world image denoising is a fundamental task in low-level vision. However, it is challenging due to the difficulty of obtaining noisy-clean image pairs in real-world scenarios. One of the most efficient approaches to address this challenge is to model the noise of imaging devices, such as smartphone cameras. This allows us to create synthetic noisy-clean image pairs by adding simulated noise to clean images, which are then used to train neural networks for image denoising. Therefore, accurate noise modeling is crucial for effective real-world image denoising. Moreover, accurate noise modeling not only benefits image denoising  [23, 27] but also improves performance in various computer vision tasks such as object detection and segmentation [24].
@@ 1 Introduction
Numerous studies have been conducted on accurate noise modeling. The simplest and most commonly used noise models include additive white Gaussian noise (AWGN) and Poisson noise. However, since the real-world noise model is highly complex, these simple noise models are insufficient to accurately capture it. To address the issue, studies using Poisson-Gaussian noise model [11] or heteroscedastic Gaussian noise model [10, 2] have been proposed. While these methods have shown improved performance over previous simple noise models, they still do not fully capture the complex distribution of real-world noise.
@@
Recently, numerous data-driven methods [1, 5, 15, 25, 7, 32] based on deep learning have been proposed. Because these methods are more sophisticated than previous methods, they are able to represent more complex distributions. They use neural networks trained on large-scale datasets of noisy images to learn the noise distribution. Specifically, they focus on modeling both signal-dependent and signal-independent noise from raw sensor images. These methods can model noise much more accurately, and the denoising neural networks trained on the modeled noise also show promising performance.
@@
As successful noise modeling has been achieved from raw sensor images, studies have been proposed that perform noise modeling in the standard RGB (sRGB) color space [20, 12, 16, 21, 32]. Generally, noise modeling is categorized based on the image domain it is conducted in, whether in the raw-RGB or the sRGB domains. Images in the raw-RGB domain are unprocessed raw sensor images without the application of image signal processing (ISP), maintaining the original data captured by the sensor. In contrast, images in the sRGB domain are applied to the ISP, resulting in standardized and enhanced visuals suitable for various displays and environments.
@@
Noise modeling is notably more difficult in the sRGB domain than in the raw-RGB domain due to the ISP’s impact on both clean signals and noise, resulting in a complex noise distribution [26]. Nevertheless, noise modeling in the sRGB domain is essential because the conversion from raw-RGB domain to sRGB domain is done by in-camera processing [20]. Often, image denoising is not applied or is insufficiently applied prior to the ISP. Consequently, effective noise modeling in the sRGB domain becomes necessary.
@@
Most deep learning-based noise modeling methods including those in the sRGB image domain employ generative models. This is because they can learn complex, high-dimensional data distributions, which is essential for accurately capturing the statistical properties of noise. There are a few efforts to model sRGB noise by using generative models such as Generative Adversarial Network (GAN) [13], Normalizing Flows [29, 28]. These methods have shown promising performance in sRGB noise modeling. However, since each of these generative models has its own strengths and weaknesses, their performance is limited.
@@
In GAN-based methods [12, 16, 21], because the characteristics of GAN are well suited to handling high-dimensional and complex data distributions, they can generate realistic noise images in the sRGB domain. However, GAN-based methods have limitations in training stability when dealing with small-sized datasets. Unfortunately, noise modeling is often conducted with insufficient datasets because generating noisy-clean pairs is a highly time-consuming task. Furthermore, since the noise distribution is affected by various camera conditions (e.g. types, settings)[2, 4], it is essential to employ a generative model that can maintain high training stability even with small-sized datasets containing limited examples of each specific condition. In contrast to GAN-based methods, Normalizing Flows-based methods [20] show higher training stability even with small-sized datasets by directly learning the probability density function of the data. However, due to the constraints of Normalizing Flows where each layer must be invertible, there are limitations on the expressiveness of the transformation function. Therefore, generating realistic noise images is constrained compared to GAN-based methods.
1.0.1 Contributions.
@@
In this paper, we propose a novel data-driven sRGB noise modeling method named NM-FlowGAN, which is a hybrid approach that combines Normalizing Flows and GAN. Our method is designed to effectively model the highly complex noise distribution found in the sRGB image domain by exploiting the strengths of each generative model. Firstly, we take advantage of the high training stability offered by Normalizing Flows to model pixel-wise noise. Our analysis of noise in the sRGB image domain reveals that pixel-wise noise depends on the clean image intensity, surrounding image structures, and camera conditions. To address this, we design novel invertible layers that can effectively model these dependencies. Furthermore, we employ GAN to model high-dimensional noise features, such as spatial correlations, which are induced by in-camera imaging processes such as demosaicing. This approach enables the generation of realistic sRGB noise images while ensuring high training stability. To validate our method, we compared the performance of noise modeling in the sRGB image domain in real-world scenarios. Our method shows superior performance compared to the other baselines.
2 Related Works
2.1 Generative Adversarial Network
@@
To address the issue, a noise modeling method NeCA [12] is proposed, which estimates the noise level using an additional noisy image and models spatially correlated noise [34] using GAN. While this method can generate realistic noise images, its performance is significantly affected by the arbitrary noisy images used in gain estimation. In addition, the method has the problem that it requires extensive training time due to the necessity of training neural networks for each camera type.
@@
2.2 Normalizing Flows
@@
To sum up, while both GAN and Normalizing Flows have been actively adopted for noise modeling in the sRGB image domain, each method has its own strengths and weaknesses. Unlike previous attempts to model noise with a generative model, we model the noise with a hybrid approach based on both models by exploiting the strengths of each model. We exploit the advantage of GAN in extracting high-dimensional noise features to effectively model spatial correlation and the strength of Normalizing Flows in maintaining high training stability even with small-sized datasets. This strength of Normalizing Flows allows our proposed neural network to effectively capture the factors affecting the noise distribution, such as camera conditions, without the necessity of training separate neural networks for each camera type.
3 Method
3.1 Analyzing sRGB Noise
sRGB noise has unpredictable and complex distributions. For accurate noise modeling, identifying the factors that affect its distribution is crucial. Therefore, we analyzed the distributions of sRGB noise from the SIDD dataset under different scenarios. Our analysis of the SIDD dataset, summarized in Figs. 1 and 2, confirms that sRGB noise is affected by three main factors: clean image intensity, image structure, and spatial correlation. For accurate noise modeling, it is essential to consider all factors that affect the noise. Detailed descriptions of each factor are given below.
@@ 3.1 Analyzing sRGB Noise
where $I(\cdot)$ denotes the ISP. From this equation, we can deduce that sRGB noise is affected by the noise present in raw-RGB; therefore, it can be affected by the clean image intensity, camera type, and ISO setting. Figs. 0(a) and 0(b) demonstrate that sRGB noise is affected by the clean image intensity in a complex, non-linear fashion and varies depending on the camera type and ISO setting.
In addition, the distribution of noise in the sRGB domain is also affected by image structures. This is evident when comparing Figs. 0(b) and 0(c); despite having the same camera type, ISO settings, and illuminant settings, the noise distribution can vary depending on the captured scene. This variability is due to signal processing techniques, such as local adjustments in the ISP. Therefore, to estimate the noise variation of a specific pixel, it is essential to consider the surrounding image structure.
Figure 2: An illustration of the correlation versus distance between neighboring pixels under various camera types and ISO settings. This figure shows that the noise value of a pixel is highly correlated with its neighboring pixels and shows similar behavior across camera types or ISO settings.
@@
where $\delta$ and $\gamma$ denote the camera type and ISO settings respectively, $P(\cdot)$ denotes the local image patch composed of the pixel $x^{\prime}_{i,c}$ and its neighboring pixels, and $f(\cdot)$ denotes the non-linear relationship between the parameters of the noise distribution and the factors that affect it. This allows us to model the pixel-wise noise $v$ at each pixel as a Gaussian distribution:
@@
The final factor to consider in the distribution of sRGB noise is spatial correlation. We have previously introduced pixel-wise noise, assuming that the distribution of sRGB noise is independent for each pixel. However, as shown in Fig. 2, the real sRGB noise of each pixel is highly correlated with the noise value of its neighboring pixels, and this relationship should be sufficiently considered during noise synthesis. This is due to the demosaicing process of the ISP on the Bayer filter [6, 17], which interpolates from neighboring noisy subpixels. It is worth noting that, in contrast to other factors that affect the distribution of pixel-wise noise, spatial correlation is unaffected by variables such as camera type or ISO settings [22]. Therefore, we can represent the sRGB noise of each pixel, considering spatial correlation, through the following equation:
@@
where $g(\cdot)$ denotes a function that maps the spatial correlation between the pixel $v_{i,c}$ of the noise image and its neighboring pixels.
3.2 Hybrid Architecture
Inspired by our analysis, we introduce our novel sRGB noise modeling method named NM-FlowGAN, which is a hybrid approach that combines Normalizing Flows and GAN. As shown in Fig. 3, our proposed neural network for synthesizing real sRGB noise includes two main components: the pixel-wise noise modeling network and the spatial correlation modeling network.
Figure 3: The overall architecture of our proposed framework for noise synthesis. Our NM-FlowGAN is comprised of two main components: the pixel-wise modeling network and the spatial correlation modeling network. These networks are based on Normalizing Flows and GAN, respectively. We employ a dequantization layer at the beginning of the pixel-wise noise modeling network, which adds uniformly sampled noise to the images during training. For better visualization, the magnitudes of the noise images are amplified.

### v3

# NM-FlowGAN: Modeling sRGB Noise without Paired Images using a Hybrid Approach of Normalizing Flows and GAN

@@ Abstract
Modeling and synthesizing real sRGB noise is crucial for various low-level vision tasks, such as building datasets for training image denoising systems. The distribution of real sRGB noise is highly complex and affected by a multitude of factors, making its accurate modeling extremely challenging. Therefore, recent studies have proposed methods that employ data-driven generative models, such as Generative Adversarial Networks (GAN) and Normalizing Flows. These studies achieve more accurate modeling of sRGB noise compared to traditional noise modeling methods. However, there are performance limitations due to the inherent characteristics of each generative model. To address this issue, we propose NM-FlowGAN, a hybrid approach that exploits the strengths of both GAN and Normalizing Flows. We combine pixel-wise noise modeling networks based on Normalizing Flows and spatial correlation modeling networks based on GAN. Specifically, the pixel-wise noise modeling network leverages the high training stability of Normalizing Flows to capture noise characteristics that are affected by a multitude of factors, and the spatial correlation networks efficiently model pixel-to-pixel relationships. In particular, unlike recent methods that rely on paired noisy images, our method synthesizes noise using clean images and factors that affect noise characteristics, such as easily obtainable parameters like camera type and ISO settings, making it applicable to various fields where obtaining noisy-clean image pairs is not feasible. In our experiments, our NM-FlowGAN outperforms other baselines in the sRGB noise synthesis task. Moreover, the denoising neural network trained with synthesized image pairs from our model shows superior performance compared to other baselines.
Index Terms:
sRGB Noise Modeling, Noise Synthesizing, Image Denoising
I Introduction
Real-world image denoising is a fundamental task in low-level vision. However, training effective neural networks for real-world image denoising is highly challenging, due to the difficulty in obtaining noisy-clean image pairs for training image denoising neural networks. A noisy-clean image pair consists of a noisy image and a corresponding clean image that shares the same noise-free content. Because noise is affected by factors such as lighting conditions, sensor limitations, or environmental influences, achieving the perfect alignment between noisy and clean images is a significant challenge.
Therefore, existing real-world image denoising datasets have been limited to those captured in controlled environments, where stationary imaging devices are used to capture static scenes [1, 2, 3]. Consequently, image denoising neural networks trained solely on data obtained under such conditions may exhibit degraded performance when processing images from uncontrolled environments, such as dynamic scenes.
(a) Training noise modeling networks
(b) Training image denoising networks
Figure 1: Two-step pipeline for training real-world image denoising networks with noise modeling networks. (a) First, our method trains noise modeling networks called NM-FlowGAN to synthesize noise $\tilde{n}$ using clean images $x$ and camera parameters $\delta$ and $\gamma$, which represent the camera type and ISO setting. (b) Once trained, NM-FlowGAN synthesizes noise that is added to clean images, generating synthesized noisy-clean image pairs that are used to train image denoising networks.
One effective approach to overcome this challenge is noise modeling, which involves estimating noise characteristics of imaging devices such as smartphone cameras. Specifically, noise modeling neural networks can be trained using data captured in controlled environments to model the noise from imaging devices. Once trained, these noise modeling neural networks can synthesize noise into clean images captured in various environments, generating a large amount of noisy-clean image pairs for training image denoising networks. This method can generate noisy-clean image pairs even in uncontrolled environments where obtaining noisy-clean image pairs is not feasible due to physical constraints. Consequently, image denoising neural networks trained on such synthesized noisy-clean image pairs can demonstrate robust performance on scenes captured in various environments. Figure 1 illustrates the process of training a noise modeling network and using it to train image denoising neural networks.
Furthermore, noise modeling can also be applied to medical imaging [4], such as X-ray imaging, where the acquisition of noisy-clean image pairs is particularly challenging. In X-ray imaging, capturing multiple images of the same scene can result in unnecessary radiation exposure, raising serious health concerns for patients. Additionally, involuntary movements of internal organs make it difficult to acquire perfectly aligned image pairs, which are essential for supervised learning in denoising tasks. By utilizing noise modeling techniques that rely solely on clean images and known imaging parameters, we can generate synthetic noisy images without subjecting patients to additional radiation or relying on impractical imaging conditions. This approach not only preserves patient safety but also facilitates the development of more effective denoising algorithms for medical applications.
@@ I Introduction
Due to these advantages, numerous studies have focused on accurate noise modeling. The simplest and most commonly used noise models include additive white Gaussian noise (AWGN) and Poisson noise. However, these simple models are insufficient to capture the complex distribution of real-world noise. To address this issue, more advanced models like the Poisson-Gaussian noise model [5] and the heteroscedastic Gaussian noise model [6, 1] have been proposed. While these methods have shown improved performance over simpler models, they still do not fully encapsulate the intricate characteristics of real-world noise.
@@
Recently, numerous data-driven methods [7, 8, 9, 10, 11, 12] based on deep learning have been proposed. These methods are more sophisticated than previous ones, capable of representing more complex distributions. They use neural networks trained on large-scale datasets of noisy images to learn the noise distribution, focusing on modeling both signal-dependent and signal-independent noise from raw sensor images. These approaches can model noise much more accurately, and the denoising neural networks trained on the modeled noise also show promising performance.
@@
Building on the success of noise modeling from raw sensor images, studies have been proposed to perform noise modeling in the standard RGB (sRGB) color space [13, 14, 15, 16, 12]. Generally, noise modeling is categorized based on the image domain in which it is conducted, either in the raw-RGB or the sRGB domains. Images in the raw-RGB domain are unprocessed raw sensor data without the application of image signal processing (ISP), maintaining the original data captured by the sensor. In contrast, images in the sRGB domain have undergone the ISP, resulting in standardized and enhanced visuals suitable for various displays and environments.
@@
Noise modeling is notably more difficult in the sRGB domain than in the raw-RGB domain due to the ISP’s impact on both clean signals and noise, resulting in a complex noise distribution [17]. Nevertheless, noise modeling in the sRGB domain is essential because the conversion from raw-RGB to sRGB is performed by in-camera processing [13]. Often, image denoising is not applied or is insufficiently applied prior to the ISP. Consequently, effective noise modeling in the sRGB domain becomes necessary.
@@
Most deep learning-based noise modeling methods in the sRGB domain employ generative models because they can learn complex, high-dimensional data distributions, which is essential for accurately capturing the statistical properties of noise. Efforts have been made to model sRGB noise using generative models such as Generative Adversarial Networks (GAN)  [18] and Normalizing Flows  [19, 20]. These methods have shown promising performance in sRGB noise modeling; however, each of these generative models has its own strengths and weaknesses, limiting their effectiveness.
@@
GAN-based methods [14, 15, 16] are well-suited to handling high-dimensional and complex data distributions, enabling them to generate realistic noise images in the sRGB domain. However, they have limitations in training stability when dealing with small-sized datasets. Noise modeling often involves insufficient datasets because generating noisy-clean pairs is highly time-consuming. Additionally, the noise distribution is affected by various camera conditions (e.g., types, settings) [1, 2]. Therefore, it is crucial to employ a generative model that can maintain high training stability even with limited data for each specific condition.
In contrast, Normalizing Flows-based methods [13, 21] exhibit higher training stability with small datasets by directly learning the probability density function of the data. However, due to the constraints of Normalizing Flows, where each layer must be invertible, there are limitations on the expressiveness of the transformation functions. Consequently, generating realistic noise images remains challenging compared to GAN-based methods.
To address the limitations of both GAN-based and Normalizing Flows-based methods, we propose a novel data-driven sRGB noise modeling method named NM-FlowGAN, which is a hybrid approach that combines the strengths of Normalizing Flows and GAN. Our method is designed to effectively model the highly complex noise distribution found in the sRGB image domain. By integrating the high training stability of Normalizing Flows with the powerful expressiveness of GAN, we can overcome the individual limitations of each method.
Moreover, unlike recently proposed methods such as NeCA-W [14] and NAFlow [21], which have demonstrated promising noise modeling performance but rely on real noisy images paired with clean images [^1] This implies that additional synthesized noisy-clean image pairs are generated from the original noisy-clean pair., our method synthesizes noisy images without such pairs. Using real paired noisy images enables the generation of realistic noise and can enhance the performance of image denoising neural networks through data augmentation. However, the requirement for paired real noisy images to synthesize noise makes it impossible to apply these methods in environments where obtaining noisy-clean image pairs is not feasible. By relying only on clean images and factors that affect noise characteristics, such as camera type or ISO settings, our method broadens the applicability of noise modeling to various environments such as dynamic scenes.
@@
In our proposed method, we first take advantage of the high training stability offered by Normalizing Flows to model pixel-wise noise. Our analysis of noise in the sRGB domain reveals that pixel-wise noise depends on the clean image intensity, surrounding image structures, and camera conditions. To address this, we design novel invertible layers that can effectively model these dependencies. Additionally, we employ GAN to model high-dimensional noise features, such as spatial correlations induced by in-camera imaging processes like demosaicing. This combined approach enables the generation of realistic sRGB noise images while ensuring high training stability. To validate our method, we compared the performance of noise modeling in the sRGB image domain in real-world scenarios. In the experiments, our method shows superior performance compared to the other baselines.
II Related Works
II-A Generative Adversarial Network
@@
To address the issue, a noise modeling method NeCA [14] is proposed, which estimates the noise level using an additional noisy image and models spatially correlated noise [22] using GAN. While this method can generate realistic noise images, its performance is significantly affected by the noisy image used in gain estimation. In addition, the method has the problem that it requires extensive training time due to the necessity of training neural networks for each camera type.
@@
(a) Scene: 8, Cam: G4, ISO: 100
(b) Scene: 8, Cam: IP, ISO: 400
(c) Scene: 7, Cam: IP, ISO: 400
II-B Normalizing Flows
@@
To address these challenges, Kim et al. [21] proposed a method called NAFlow, which focuses on improving the handling of spatially correlated noise through the use of multi-scale noise embedding techniques and Gaussian mixture models. By incorporating these techniques, NAFlow effectively models the spatial correlation of noise. However, NAFlow has a limitation in that it uses real-world noisy images paired with clean images to synthesize the noise. Using real noisy-clean image pairs limits NAFlow’s ability to perform noise synthesis in real-world scenarios. This constraint specifically restricts its applicability, such as when processing images captured in uncontrolled environments such as dynamic scenes where obtaining noisy-clean image pairs is not feasible. As a result, the challenge of effectively modeling the spatial correlation of noise using Normalizing Flows remains.
To sum up, while both GAN and Normalizing Flows have been actively adopted for noise modeling in the sRGB image domain, each method has its own strengths and weaknesses. Unlike previous attempts to model noise with a single generative model, we propose a hybrid approach that leverages the strengths of both models. By exploiting GAN’s capability in extracting high-dimensional noise features to model spatial correlation and Normalizing Flows’ high training stability with small-sized datasets, our approach effectively captures the factors affecting noise distribution, such as camera conditions, without requiring separate neural networks for each camera type. Additionally, our method does not require the paired noisy image corresponding to the clean image for noise synthesis, further enhancing its applicability in diverse real-world scenarios.
III Method
Figure 3: An illustration of the correlation $r$ versus distance $d$ between neighboring pixels under various camera types and ISO settings. This figure shows that the noise value of a pixel is highly correlated with its neighboring pixels and shows similar behavior across camera types or ISO settings.
III-A Analyzing sRGB Noise
sRGB noise has unpredictable and complex distributions. For accurate noise modeling, identifying the factors that affect its distribution is crucial. Therefore, we analyzed the distributions of sRGB noise from the SIDD dataset under different scenarios. Our analysis of the SIDD dataset, summarized in Figure 2 and 3, confirms that sRGB noise is affected by three main factors: clean image intensity, image structure, and spatial correlation. For accurate noise modeling, it is essential to consider all factors that affect the noise. Detailed descriptions of each factor are given below.
@@ III-A Analyzing sRGB Noise
where $I(\cdot)$ denotes the ISP. From this equation, we can deduce that sRGB noise is affected by the noise present in raw-RGB; therefore, it can be affected by the clean image intensity, camera type, and ISO setting. Figure 2a and 2b demonstrate that sRGB noise is affected by the clean image intensity in a complex, non-linear fashion and varies depending on the camera type and ISO setting.
In addition, the distribution of noise in the sRGB domain is also affected by image structures. This is evident when comparing Figure 2b and 2c; despite having the same camera type, ISO settings, and illuminant settings, the noise distribution can vary depending on the captured scene. This variability is due to signal processing techniques, such as local adjustments in the ISP. Therefore, to estimate the noise variation of a specific pixel, it is essential to consider the surrounding image structure.
@@
where $\delta$ and $\gamma$ denote the camera type and ISO settings respectively, $P(\cdot)$ denotes the local image patch composed of the pixel $x_{i,c}$ and its neighboring pixels, and $f(\cdot)$ denotes the non-linear relationship between the parameters of the noise distribution and the factors that affect it. This allows us to model the pixel-wise noise $v$ at each pixel as a Gaussian distribution:
@@
The final factor to consider in the distribution of sRGB noise is spatial correlation. Previously, we introduced pixel-wise noise, assuming that the distribution of sRGB noise is independent for each pixel. However, as shown in Figure 3, the actual sRGB noise of each pixel is highly correlated with the noise values of its neighboring pixels. This spatial correlation arises due to the demosaicing process of the ISP on the Bayer filter [23, 24], which interpolates from neighboring noisy subpixels. Therefore, it is crucial to account for this correlation during noise synthesis.
Spatial correlation arises mainly due to the Bayer demosaicing process, which is typically performed without regard to camera settings [25, 26, 27]. As shown in Figure 3, spatial correlation is consistent regardless of camera type or ISO settings, unlike pixel-wise noise distribution, which is affected by a multitude of factors.
To analyze this inter-pixel relationship shown in Figure 3, we used Pearson correlation. The correlation $r$ is calculated as:
$r=Cov(N,\tilde{N})/(\sigma_{N}\cdot\sigma_{\tilde{N}}),$ (5)
where $N$ represents the set of pixels in the noise image, and $\tilde{N}$ denotes the set of pixels at a specific distance from each pixel:
$\tilde{N}(x,y)=N(x+d_{x},y+d_{y}),$ (6)
with the distance $d$ between pixels calculated as:
$d=\sqrt{d_{x}^{2}+d_{y}^{2}}.$ (7)
Considering spatial correlation, the sRGB noise for each pixel can be modeled using the following equation:
@@
where $g(\cdot)$ denotes a function that maps spatial correlation between the pixel $v_{i,c}$ of the noise image and its neighboring pixels.
III-B Hybrid Architecture
Inspired by our analysis, we introduce our novel sRGB noise modeling method named NM-FlowGAN, which is a hybrid approach that combines Normalizing Flows and GAN. As shown in Figure 4, our proposed neural network for synthesizing real sRGB noise includes two main components: the pixel-wise noise modeling network and the spatial correlation modeling network.
Figure 4: The overall architecture of our proposed framework for noise synthesis. Our NM-FlowGAN is comprised of two main components: the pixel-wise noise modeling network and the spatial correlation modeling network. These networks are based on Normalizing Flows and GAN, respectively. We employ a dequantization layer at the beginning of the pixel-wise noise modeling network, which adds uniformly sampled noise to the images during training. For better visualization, the magnitudes of the noise images are amplified.